OpenStudy (anonymous):

Kevin and Randy Muise have a jar containing 80 coins, all of which are either quarters or nickels. The total value of the coins in the jar is $10.80. How many of each type of coin do they have?

OpenStudy (welshfella):

a nickel and a quarter I meant

OpenStudy (anonymous):

nickel: 5 cent and Dime is 10 cent

OpenStudy (welshfella):

and a quarter is 25 cents?

OpenStudy (welshfella):

you create 2 equations let x be the number of nickels and y the number of quarters so one equation is x + y = 80 (because there are a total of 80 coins) now consider the values of the coins 5x + 25y = 1080 (10 dollr 80 = 1080 cents)

OpenStudy (welshfella):

so solve for x + y: x + y = 80 5x + 25y = 1080

OpenStudy (welshfella):

(if i'm right about the quarter)